# Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge

### v3

@@ Abstract.
Language Models (LMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance diminishes when dealing with less-popular or low-frequency concepts and entities, for example in domain specific applications. The two prominent approaches to enhance the performance of LMs on low-frequent topics are: Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LMs in handling low-frequency entities on question answering tasks. We conduct extensive experiments on twelve LMs of varying size and type and different fine tuning, data augmentation, and retrieval models. Our findings indicate that while FT boosts the performance across entities of varying popularity, RAG surpasses FT by a large margin particularly for least popular factual knowledge. Additionally, the success of both RAG and FT approaches is amplified by improving retrieval and data augmentation techniques. Fine tuning, while beneficial for small LMs, requires extensive resources. To address this issue, we propose the new Stimulus RAG approach that surpasses the effectiveness of fine tuning based approaches, thereby eliminating the need for the costly data augmentation and fine tuning step for enriching LMs with less popular factual knowledge.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 2nd International ACM SIGIR Conference on Information Retrieval in the Asia Pacific 2024; December 9-12, 2024; Tokyo, Japanccs: Information systems Question answeringccs: Information systems Novelty in information retrievalccs: Information systems Language models
@@ 1. Introduction
Language Models (LMs) exhibit outstanding capabilities in executing tasks that demand extensive memorization of factual data (Chowdhery et al., 2023). However, their memorization capabilities are constrained when dealing with less frequent entities (Mallen et al., 2023; Gerritse et al., 2022; Kandpal et al., 2023; Sun et al., 2023), and even the largest models may encounter the well-known ”hallucination” problem (Shuster et al., 2021) and temporal degradation (Kasai et al., 2022). Consequently, when LMs are intended for deployment in less resourced domains, customization becomes imperative to ensure optimal performance. A common example is within the industrial setup, where chatbots or Question Answering (QA) systems need to accurately answer users’ questions about a proprietary knowledge graph or intra-company terminology with limited textual description (Zhang et al., 2024a; Soudani et al., 2023).
@@
Retrieval-Augmented Generation (RAG) and Fine-Tuning (FT) stand out as two prominent approaches for adapting LMs to specific domains (Soudani et al., 2023; Mosbach et al., 2023; de Luis Balaguer et al., 2024; Ovadia et al., 2023). RAG retrieves relevant information from a document corpus and enhances LM’s response generation through the implementation of in-context learning (ICL) (Zhang et al., 2024a; Cuconasu et al., 2024). Conversely, FT approach updates model weights to become adept at recalling specific information and enhance its memorization capabilities during inference (Asai et al., 2024). In the context of less popular knowledge, where limited data is available, data augmentation methods are utilized to generate synthetic training data, serving as an initial step towards FT (Soudani et al., 2024b, a).
@@
In this paper, we aim to understand which approach is more appropriate for customizing LMs for less resourced domains. Specifically, we seek to answer this research question: (RQ1) How RAG compares to fine-tuning for question answering over less popular factual knowledge and which factors affect their performance? To address this question, we conduct a comprehensive comparison of RAG and fine-tuning methods for less popular knowledge, assuming that textual descriptions, albeit limited, are available for a specific domain and application. We therefore collect Wikipedia documents related to QA datasets over long-tail entities and apply two methods of knowledge injections: parametric knowledge injection using FT and non-parametric knowledge injection using RAG. For the FT approach, the LM is fine-tuned with synthetically generated QAs from these documents using data augmentation approaches (Alberti et al., 2019; Ushio et al., 2023). For the RAG approach, we use retrievers to rank the most relevant documents for a query. We investigate how the effectiveness of these methods is affected by the following aspects: (i) fine-tuning method; i.e., full FT vs. parameter efficient fine-tuning (PEFT), (ii) data augmentation method, (iii) LM type and size; i.e., decoder only vs. encoder-decoder models and varying size, ranging from 80M to 11B parameters, and (iv) retrieval model performance. Through exhaustive experimentation on twelve LMs and different setup of fine-tuning, data augmentation, and retrieval models, we arrive to the following conclusions:
@@
  • Fine-tuning method: Comparing full FT with PEFT (i.e., QLoRA (Dettmers et al., 2023)) for LMs with less than 2 billion parameters, full FT is more effective in the downstream task performance. PEFT, however, preserves the reasoning ability of LMs needed for RAG and outperforms full FT when the fine-tuned models used in combination with RAG (Table 2).
@@
  • Data augmentation method: Comparing prompt-based QA generation with QA generation by a state-of-the-art fine-tuned model (Ushio et al., 2023), the prompt-based QA generation demonstrates better performance for the factual QA task. This suggests that large LMs have a superior ability to generate high-quality QA pairs, although the fine-tuned model produce a greater volume of data (Table 2).
@@
  • LM type and size: Comparing decoder-only with encoder-decoder LMs (Flan-T5 models of various sizes), decoder-only models outperform encoder-decoder models of similar size. Interestingly, larger LMs generally do not benefit from fine-tuning, while smaller do. Therefore, a small fine-tuned LM with RAG can preform on par or better than a large LM; e.g., StableLM2 (1.6B) vs. Llama3 (8B) (Table 3).
@@
  • Fine-tuning vs. RAG: Comparing these two knowledge injection methods, RAG substantially outperforms fine-tuning. with an ideal retriever, fine-tuned LMs with RAG models either outperform or perform on par with vanilla LMs with RAG in all but one case (Figure 1).
While fine-tuning improves accuracy in answering factual questions, both with and without RAG, it demands a considerable amount of effort and resources. This leads us to our second research question: (RQ2): Can we avoid the cost of fine-tuning by developing an advanced RAG approach that surpass the performance of fine-tuned LM with RAG? To answer this, we develop Stimulus RAG (SRAG), a new RAG approach that stimulates an LM to generate the correct response based on the provided hint in the prompt. The hint is extracted from the top retrieved documents by the retrieval model. Our results demonstrate that Stimuls RAG outperforms all other combinations of fine-tuning, both with and without retrieve-the-generate RAG.
@@
  • We study the effectiveness of fine-tuning and RAG approaches for question answering over less popular factual knowledge and compare the performance of these models across distinct setups: vanilla and fine-tuned models, both with and without RAG.
  • We perform extensive experiments to understand how FT and RAG models are affected by four different factors: data augmentation method, fine-tuning method, LM type and size, and retrieval model.
@@ 2. Related Work
Parametric and Non-parametric Knowledge. It is demonstrated that large pre-trained LMs memorize the significant amount of world knowledge in their parameters (parametric knowledge(Mallen et al., 2023). FT can update the parametric knowledge embedded in LMs and customize it for specific domain (Mosbach et al., 2023; Gerritse et al., 2022). One of the important principle for FT is data availability, which is limited specifically in specialized domains (de Luis Balaguer et al., 2024; Ovadia et al., 2023). Data augmentation (DA) addresses data scarcity problem by generating task- and domain-relevant samples from existing unlabeled texts. A common DA approach for the QA task is generating question-answer pairs through a four-step Pipeline, consisting of: passage selection, answer extraction, question generation, and consistency filtering (Alberti et al., 2019; Lewis et al., 2019, 2021; Ushio et al., 2023). Ushio et al. (2023) conducted an empirical study comparing three QA generation approaches: Pipeline, Multitask, and End-to-End (E2E) and showed the E2E approach outperforms others in downstream tasks.
Furthermore, a large body of work shows that augmenting LMs with nonparametric knowledge (i.e., retrieved text chunks) enables much smaller models to match the performance of larger models (Lewis et al., 2020). In this method, known as Retrieval Augmented Generation (RAG), an information retrieval system is utilized to find relevant documents and adds them to the input prompt to enhance response generation of LMs (Asai et al., 2023a, b).
@@
RAG Development. RAG introduces a new approach in AI, combining the strengths of both retrieval-based and generative models (Cuconasu et al., 2024). The concept of RAG was coined and popularized by Lewis et al. (2020), who introduced a model that combines a dense passage retriever with a sequence-to-sequence model, called Retrieve-then-Generate. This approach demonstrated substantial improvements in knowledge-intensive tasks. Several parameters affect a RAG system’s accuracy, including the performance of the retriever model (Cheng et al., 2023), the relevance of the passages included in the prompt context, their position, and their number (Cuconasu et al., 2024; Asai et al., 2024).
@@ 3.1. Task Definition
This study specifically focuses on factual knowledge (Adams, 2015) of entities, defined as information about particular attributes and characteristics of target entities, among various types of world knowledge (Mallen et al., 2023). We chose this because the amount of knowledge memorized by LMs can be approximated by its accuracy in answering simple factual questions, such as ”In what city was Lisa Miller born?” (Sun et al., 2023).
Factual knowledge is defined as a triplet of (subject, relationship, object) (Mallen et al., 2023). In this context, the question involves the subject and the relationship, while the answer corresponds to the object. By using this template questions, we ensure that LMs understand the question and derive the answer from their embedded knowledge. We select Wikipedia-based question-answering datasets focused on factual knowledge. This enable us to measure the popularity of entities based on Wikipedia pageviews and also obtain the corresponding evidence document for each entity from Wikipedia. For each dataset, we select Wikipedia pages whose corresponding entities appear in the test dataset. This setup mirrors real-world industry practices, where entities and their textual descriptions relate to companies’ specific internal concepts.
@@ 3.2. Knowledge Injection with Fine-tuning
where $q_{i}$ and $a_{i}$ denote a question and answer generated for the document $d_{i}$, $m$ is the total number of generated QAs for the document, and the generated set of question-answer pairs is denoted as $Q_{i}$. The QA generation method $\mathcal{G}_{qa}$ is applied to all documents $D$, and the generated set $Q=\bigcup_{i=1}^{n}Q_{i}$ is then used for fine tuning an LLM, which consists of a set of learnable parameters to predict the probabilities of future or masked tokens.
@@ 3.3. Knowledge Injection with RAG
The non-parametric knowledge injection is performed using RAG, which consists of two components: the Retriever and the Generator (Cuconasu et al., 2024; Asai et al., 2024).
@@ 3.4. Stimulus RAG
While the generic retrieve-then-generate framework of RAG is effective in answering factual knowledge (Mallen et al., 2023), it sometimes struggles to accurately respond factual questions that are not memorized in LMs parameters, even when the ground truth document is included in the prompt. We hypothesize that by employing a more advanced RAG approach we can achieve or even surpass the benefits of the combined RAG and FT (+FT+RAG) setup. This would allow us to bypass the computationally expensive process of knowledge injection with FT, which involves resource-intensive processes of QA generation and FT.
We introduce Stimulus RAG, a RAG approach that guides LMs to generate responses using a hint provided in the RAG prompt. Stimulus RAG comprises three steps: Retriever, Hint extractor, and Generator. The Retriever and Generator are as defined in Section 3.3. The Hint extractor provide hint text that guides LM to generate accurate responses and works as follows: The top-$K$ ranked documents $D_{q}^{K}=[d_{1},...,d_{K}]$ ($d_{K}$ denoting the retrieved document at rank $K$) from the retriever step are split into sentences, denoted as $S_{q}=\{s_{j}\}_{j=0}^{N}$, where $N$ is the total number of sentences. These sentences are then ranked using a retrieval model $R^{\prime}(S_{q},q)$. For the sake of simplicity, we use the same retrieval model as that used in the retriever step, hence $R=R^{\prime}$. The top-ranked sentence $s_{q}^{1}$ is identified as the most relevant sentence for the given question and serves as a hint. The sentence $s_{q}^{1}$ can be directly used in the prompt, which we refer to as the SRAG(S) approach. Alternatively, the document containing the sentence $s_{q}^{1}$ can be considered as a hint, which we refer to as the SRAG(D) approach. The extracted hint is placed at the top of the input prompt provided to the Generator component. This implies that the hint text a repetition of presumably the most relevant sentence/document of in the prompt. It has been shown that the beginning of the input prompt receives more attention from LMs (Liu et al., 2023). Our method is illustrated in Figure 4.
@@ 4. Experimental Setup
Evaluation Metric. Following previous studies (Mallen et al., 2023; Maekawa et al., 2024; Sciavolino et al., 2021), we report on the accuracy metric, where a prediction is considered correct if one of the ground truth responses matches a substring of the predicted response. While widely used (Kandpal et al., 2023; Liu et al., 2023; Cuconasu et al., 2024), this metric is not without issues. A principal problem arises in determining response correctness, particularly in cases involving date representations or varying phrasings that convey identical meanings (Cuconasu et al., 2024). For example, comparing the predicted response “Nathanson” with ground truth “Jeff Nathanson,” the prediction is considered incorrect. Another observed problem is that when the model generates multiple entity names, among them the ground truth entity, the response is incorrectly considered as correct. Recognizing these limitations, we acknowledge the necessity for a more advanced analysis of answer variations, which we leave for future research.
@@
Language Models. We use several LMs, focusing on two main features: the backbone architecture (i.e., decoder-only and encoder-decoder) and model size, which ranges from 80 million to over 11 billion parameters. For the encoder-decoder models, we utilize five versions of the FlanT5 family (Chung et al., 2022), spanning from small to XXL. For the decoder-only models, we employ smaller LMs such as Tiny-llama (Zhang et al., 2024b), StableLM2 (Bellagente et al., 2024), and MiniCPM (Hu et al., 2024), which range from 1 to 2 billion parameters. Additionally, we incorporate larger LMs like Mistral (Jiang et al., 2023), Zephyr (Tunstall et al., 2023), LLama2 (Touvron et al., 2023), and LLama3, [^1] https://ai.meta.com/blog/meta-llama-3 which range from 7 to 8 billion parameters. As for instructions, we apply zero-shot prompting for generative prediction using a straightforward templates. For non retrieved-augmented input, we use template: "Question: <question>", and for retrieved-augmented input, we use template: "Context: <context> Question: <question>".
@@
Fine Tuning. We generate training data for the FT approach using two distinct data augmentation methods. To ensure a fair comparison between FT and RAG with an ideal retriever, we generate QAs exclusively using the summary sections of Wikipedia pages. After generating QA pairs, we proceed to fine-tune LMs using two approaches: full parameter tuning (Full) and Parameter Efficient Fine-Tuning (PEFT). Within the range of PEFT techniques (Zaken et al., 2022; Liu et al., 2022; Ma et al., 2023), we utilize QLoRA (Dettmers et al., 2023), chosen for its broad acceptance in the field (Kaddour et al., 2023; Naveed et al., 2023) and efficient use of computational resources we had at our disposal.
@@
RAG. We utilize a variety of retrieval models to obtain relevant documents for the RAG approach, including BM25 (Robertson and Zaragoza, 2009), Contriever (Izacard et al., 2022), DPR (Karpukhin et al., 2020), and a two-stage re-ranker that combines BM25 with DPR, all implemented according to the BEIR benchmark (Thakur et al., 2021). Additionally, since the selected datasets do not contain grounded document evidence, we assume that the summary section of each Wikipedia page is the answer-containing document. We define an ideal retriever model as one that returns the summary paragraph as the top-ranked document, referred to as the Ideal retriever throughout the paper. We acknowledge that this assumption is not entirely accurate as some answers may be found in other subsections. However, our evaluation on the downstream task (Figures 6 and 8) demonstrates that the Ideal retriever outperforms other retrievers.
@@
Stimulus RAG. We select a DPR retriever for the hint extractor and set $K=3$. We report on two variation of the SRAG approach: (i) SRAG(S), which utilizes the top-1 sentence as the hint, and (ii) SRAG(D), which inserts the entire document contacting the top-ranked sentence. For the instruction, we used the following template: "Context: <hint><context> Question: <question>".

### v5

@@ Abstract.
Language Models (LMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance of these models diminishes when dealing with less-popular or low-frequency concepts, for example, in domain-specific applications. The two prominent approaches to enhance the performance of LMs on less frequent topics are Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LMs in handling low-frequency entities in question answering tasks. We conduct extensive experiments on twelve LMs of varying size and type and different FT methods, data augmentation, and retrieval models. Our findings indicate that while FT boosts performance across entities of varying popularity, RAG surpasses FT by a large margin, particularly for least popular factual knowledge. Additionally, the success of both RAG and FT approaches is amplified by improving retrieval and data augmentation techniques. Fine-tuning, while beneficial for small LMs, requires extensive resources. To address this issue, we propose the new Stimulus RAG approach that surpasses the effectiveness of fine-tuning based approaches, thereby eliminating the need for the costly data augmentation and fine-tuning steps for enriching LMs with less popular factual knowledge. The code is available at https://github.com/informagi/RAGvsFT.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 9–12, 2024; Tokyo, Japanbooktitle: Proceedings of the 2024 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP ’24), December 9–12, 2024, Tokyo, Japandoi: 10.1145/3673791.3698415isbn: 979-8-4007-0724-7/24/12ccs: Computing methodologies Natural language generationccs: Information systems Question answeringccs: Information systems Novelty in information retrievalccs: Information systems Language models
@@ 1. Introduction
Language Models (LMs) exhibit outstanding capabilities in executing tasks that demand extensive memorization of factual data (Chowdhery et al., 2023). However, their memorization capabilities are constrained when dealing with less frequent entities (Mallen et al., 2023; Gerritse et al., 2022; Kandpal et al., 2023; Sun et al., 2024), and even the largest models may encounter the well-known ”hallucination” problem (Shuster et al., 2021) and temporal degradation (Kasai et al., 2023). Consequently, when LMs are intended for deployment in less resourced domains, customization becomes imperative to ensure optimal performance. A common example is within the industrial setup, where chatbots or Question Answering (QA) systems need to accurately answer users’ questions about a proprietary knowledge graph or intra-company terminology with limited textual description (Zhang et al., 2024a; Soudani et al., 2023).
@@
Retrieval-Augmented Generation (RAG) and Fine-Tuning (FT) stand out as two prominent approaches for adapting LMs to specific domains (Soudani et al., 2023; Mosbach et al., 2023; de Luis Balaguer et al., 2024; Ovadia et al., 2023). RAG retrieves relevant information from a document corpus and enhances LM’s response generation through the implementation of in-context learning (ICL) (Zhang et al., 2024a; Cuconasu et al., 2024). Conversely, FT approach updates model weights to become adept at recalling specific information and enhance its memorization capabilities during inference (Asai et al., 2024b). In the context of less popular knowledge, where limited data is available, data augmentation methods are utilized to generate synthetic training data, serving as an initial step towards FT (Soudani et al., 2024b, a). Despite existing research on enhancing LM’s memorization with RAG (Mallen et al., 2023), no work to our knowledge has compared RAG with with knowledge obtained through FT, particularly for less popular knowledge.
@@
In this paper, we aim to understand which approach is more appropriate for customizing LMs for less resourced domains. Specifically, we seek to answer this research question: (RQ1) How does RAG compare to fine-tuning for question answering over less popular factual knowledge, and which factors affect their performance? To address this question, we conduct a comprehensive comparison of RAG and fine-tuning methods for less popular knowledge, assuming that textual descriptions, albeit limited, are available for a specific domain and application. We, therefore, collect Wikipedia documents related to QA datasets over long-tail entities and apply two methods of knowledge injection: parametric knowledge injection using FT and non-parametric knowledge injection using RAG. For the FT approach, the LM is fine-tuned with synthetically generated QAs from these documents using data augmentation approaches (Alberti et al., 2019; Ushio et al., 2023). For the RAG approach, we use retrievers to rank the most relevant documents for a query. We investigate how the effectiveness of these methods is affected by the following aspects: (i) fine-tuning method; i.e., full FT vs. parameter efficient fine-tuning (PEFT), (ii) data augmentation method, (iii) LM type and size; i.e., decoder only vs. encoder-decoder models and varying size, ranging from 80M to 11B parameters, and (iv) retrieval model performance. Through exhaustive experimentation on twelve LMs and different setup of fine-tuning, data augmentation, and retrieval models, we arrive at the following conclusions:
@@
  • Fine-tuning method: Comparing full FT with PEFT (i.e., QLoRA (Dettmers et al., 2023)) for LMs with less than 2 billion parameters, full FT is more effective than PEFT in the downstream task. PEFT, however, preserves the reasoning ability of LMs (needed for RAG) and outperforms full FT when the fine-tuned models are used in combination with RAG (Table 2).
@@
  • Data augmentation method: Comparing prompt-based and a state-of-the-art fine-tuned (Ushio et al., 2023) QA generation models, the prompt-based method demonstrates better performance for the downstream task. This suggests that the high-quality synthetic data generated by large LMs can better assist LMs with memorizing new knowledge, compared to the greater volume of data generated by the fine-tuned model (Table 2).
@@
  • LM type and size: Comparing decoder-only with encoder-decoder LMs (Flan-T5 models of various sizes), decoder-only models outperform encoder-decoder models of similar size. Interestingly, larger LMs generally do not benefit from fine-tuning, while smaller ones do. Therefore, a small fine-tuned LM with RAG can perform on par or better than a large LM; e.g., StableLM2 (1.6B) vs. Llama3 (8B) (Table 3).
@@
  • Fine-tuning vs. RAG: Comparing these two knowledge injection methods, RAG substantially outperforms fine-tuning. Fine-tuned LMs combined with RAG either outperform or perform on par with vanilla LMs with RAG in all but one case (Figure 1).
While fine-tuning improves accuracy in answering factual questions, both with and without RAG, it demands a considerable amount of effort and resources. This leads us to our second research question: (RQ2): Can we avoid the cost of fine-tuning by developing an advanced RAG approach that surpass the performance of a fine-tuned LM with RAG? To answer this question, we develop Stimulus RAG (SRAG), a new RAG approach that stimulates an LM to generate the correct response based on the provided hint in the prompt. The hint is extracted from the top retrieved documents by the retrieval model. Our results demonstrate that Stimulus RAG outperforms all other combinations of fine-tuning, both with and without retrieve-then-generate RAG.
@@
  • We study the effectiveness of fine-tuning and RAG approaches for question answering over less popular factual knowledge and compare the performance of these models across distinct setups: vanilla and fine-tuned models, both with and without RAG, using different data augmentation methods.
  • We perform extensive experiments to understand how fine-tuned and RAG models are affected by four different factors: data augmentation method, fine-tuning method, LM type and size, and retrieval model.
@@ 2. Related Work
Parametric and Non-parametric Knowledge. It is demonstrated that large pre-trained LMs memorize a significant amount of world knowledge in their parameters (parametric knowledge(Mallen et al., 2023). FT can update the parametric knowledge embedded in LMs and customize it for a specific domain (Mosbach et al., 2023; Gerritse et al., 2022). One of the important principles for FT is data availability, which is limited specifically in specialized domains (de Luis Balaguer et al., 2024; Ovadia et al., 2023). Data augmentation (DA) addresses the data scarcity problem by generating task- and domain-relevant samples from existing unlabeled texts. A common DA approach for the QA task is generating question-answer pairs through a four-step Pipeline, consisting of passage selection, answer extraction, question generation, and consistency filtering (Alberti et al., 2019; Lewis et al., 2019, 2021; Ushio et al., 2023). Ushio et al. (2023) conducted an empirical study comparing three QA generation approaches: Pipeline, Multitask, and End-to-End (E2E) and showed the E2E approach outperforms others in downstream tasks.
Furthermore, a large body of work shows that augmenting LMs with nonparametric knowledge (i.e., retrieved text chunks) enables much smaller models to match the performance of larger models (Lewis et al., 2020). In this method, known as Retrieval Augmented Generation (RAG), an information retrieval system is utilized to find relevant documents and adds them to the input prompt to enhance response generation of LMs (Asai et al., 2023, 2024a).
@@
RAG Development. RAG introduces a new approach in AI, combining the strengths of both retrieval-based and generative models (Cuconasu et al., 2024). The concept of RAG was coined and popularized by Lewis et al. (2020), who introduced a model that combines a dense passage retriever with a sequence-to-sequence model, called Retrieve-then-Generate. This approach demonstrated substantial improvements in knowledge-intensive tasks. Several parameters affect a RAG system’s accuracy, including the performance of the retriever model (Cheng et al., 2023), the relevance of the passages included in the prompt context, their position, and their number (Cuconasu et al., 2024; Asai et al., 2024b).
@@ 3.1. Task Definition
This study specifically focuses on factual knowledge (Adams, 2015) of entities, defined as information about particular attributes and characteristics of target entities, among various types of world knowledge (Mallen et al., 2023). We chose this because the amount of knowledge memorized by LMs can be approximated by its accuracy in answering simple factual questions, such as ”In what city was Lisa Miller born?” (Sun et al., 2024).
Factual knowledge is defined as a triplet of (subject, relationship, object) (Mallen et al., 2023). In this context, the question involves the subject and the relationship, while the answer corresponds to the object. By using these template questions, we ensure that LMs understand the question and derive the answer from their embedded knowledge. We select Wikipedia-based question-answering datasets focused on factual knowledge. This enables us to measure the popularity of entities based on Wikipedia pageviews and also obtain the corresponding evidence document for each entity from Wikipedia. For each dataset, we select Wikipedia pages whose corresponding entities appear in the test dataset. This setup mirrors real-world industry practices, where entities and their textual descriptions relate to companies’ specific internal concepts.
@@ 3.2. Knowledge Injection with Fine-tuning
where $q_{i}$ and $a_{i}$ denote a question and answer generated for the document $d_{i}$, $m$ is the total number of generated QAs for the document, and the generated set of question-answer pairs is denoted as $Q_{i}$. The QA generation method $\mathcal{G}_{qa}$ is applied to all documents $D$, and the generated set $Q=\bigcup_{i=1}^{n}Q_{i}$ is then used for fine-tuning an LLM, which consists of a set of learnable parameters to predict the probabilities of future or masked tokens.
@@ 3.3. Knowledge Injection with RAG
The non-parametric knowledge injection is performed using RAG, which consists of two components: the Retriever and the Generator (Cuconasu et al., 2024; Asai et al., 2024b).
@@ 3.4. Stimulus RAG
While the generic retrieve-then-generate framework of RAG is effective in answering factual knowledge (Mallen et al., 2023), it sometimes struggles to accurately respond to factual questions that are not memorized in LMs parameters, even when the ground truth document is included in the prompt. We hypothesize that by employing a more advanced RAG approach we can achieve or even surpass the benefits of the combined RAG and FT (+FT+RAG) setup. This would allow us to bypass the computationally expensive process of knowledge injection with FT, which involves resource-intensive processes of QA generation and FT.
We introduce Stimulus RAG, a RAG approach that guides LMs to generate responses using a hint provided in the RAG prompt. Stimulus RAG comprises three steps: Retriever, Hint extractor, and Generator. The Retriever and Generator are as defined in Section 3.3. The Hint extractor provide hint text that guides LM to generate accurate responses and works as follows: The top-$K$ ranked documents $D_{q}^{K}=[d_{1},...,d_{K}]$ ($d_{K}$ denoting the retrieved document at rank $K$) from the retriever step are split into sentences, denoted as $S_{q}=\{s_{j}\}_{j=0}^{N}$, where $N$ is the total number of sentences. These sentences are then ranked using a retrieval model $R^{\prime}(S_{q},q)$. For the sake of simplicity, we use the same retrieval model as that used in the retriever step, hence $R=R^{\prime}$. The top-ranked sentence $s_{q}^{1}$ is identified as the most relevant sentence for the given question and serves as a hint. The sentence $s_{q}^{1}$ can be directly used in the prompt, which we refer to as the SRAG(S) approach. Alternatively, the document containing the sentence $s_{q}^{1}$ can be considered as a hint, which we refer to as the SRAG(D) approach. The extracted hint is placed at the top of the input prompt provided to the Generator component. This implies that the hint text a repetition of presumably the most relevant sentence/document of in the prompt. It has been shown that the beginning of the input prompt receives more attention from LMs (Liu et al., 2024). Our method is illustrated in Figure 4.
@@ 4. Experimental Setup
Evaluation Metric. Following previous studies (Mallen et al., 2023; Maekawa et al., 2024; Sciavolino et al., 2021), we report on the accuracy metric, where a prediction is considered correct if one of the ground truth responses matches a substring of the predicted response. While widely used (Kandpal et al., 2023; Liu et al., 2024; Cuconasu et al., 2024), this metric is not without issues. A principal problem arises in determining response correctness, particularly in cases involving date representations or varying phrasings that convey identical meanings (Cuconasu et al., 2024). For example, comparing the predicted response “Nathanson” with the ground truth “Jeff Nathanson,” the prediction is considered incorrect. Another observed problem is that when the model generates multiple entity names, among them the ground truth entity, the response is incorrectly considered as correct. Recognizing these limitations, we acknowledge the necessity for a more advanced analysis of answer variations, which we leave for future research.
@@
Language Models. We use several LMs, focusing on two main features: the backbone architecture (i.e., decoder-only and encoder-decoder) and model size, which ranges from 80 million to over 11 billion parameters. For the encoder-decoder models, we utilize five versions of the FlanT5 family (Chung et al., 2024), spanning from small to XXL. For the decoder-only models, we employ smaller LMs such as Tiny-llama (Zhang et al., 2024b), StableLM2 (Bellagente et al., 2024), and MiniCPM (Hu et al., 2024), which range from 1 to 2 billion parameters. Additionally, we incorporate larger LMs like Mistral (Jiang et al., 2023), Zephyr (Tunstall et al., 2023), LLama2 (Touvron et al., 2023), and LLama3, [^1] https://ai.meta.com/blog/meta-llama-3 which range from 7 to 8 billion parameters. As for instructions, we apply zero-shot prompting for generative prediction using a straightforward template. For non retrieved-augmented input, we use template: "Question: <question>", and for retrieved-augmented input, we use template: "Context: <context> Question: <question>".
@@
Fine Tuning. We generate training data for the FT approach using two distinct data augmentation methods. To ensure a fair comparison between FT and RAG with an ideal retriever, we generate QAs exclusively using the summary sections of Wikipedia pages. After generating QA pairs, we proceed to fine-tune LMs using two approaches: full parameter tuning (Full) and Parameter Efficient Fine-Tuning (PEFT). Within the range of PEFT techniques (Zaken et al., 2022; Liu et al., 2022; Ma et al., 2024), we utilize QLoRA (Dettmers et al., 2023), chosen for its broad acceptance in the field (Kaddour et al., 2023; Naveed et al., 2023) and efficient use of computational resources we had at our disposal.
@@
RAG. We utilize a variety of retrieval models to obtain relevant documents for the RAG approach, including BM25 (Robertson and Zaragoza, 2009), Contriever (Izacard et al., 2022), DPR (Karpukhin et al., 2020), and a two-stage re-ranker that combines BM25 with DPR, all implemented according to the BEIR benchmark (Thakur et al., 2021). Additionally, since the selected datasets do not contain grounded document evidence, we assume that the summary section of each Wikipedia page is the answer-containing document. We define an ideal retriever model as one that returns the summary paragraph as the top-ranked document, referred to as the Ideal retriever throughout the paper. We acknowledge that this assumption is not entirely accurate as some answers may be found in other subsections. However, our evaluation of the downstream task (Figures 6 and 8) demonstrates that the Ideal retriever outperforms other retrievers.
@@
Stimulus RAG. We select a DPR retriever for the hint extractor and set $K=3$. We report on two variations of the SRAG approach: (i) SRAG(S), which utilizes the top-1 sentence as the hint, and (ii) SRAG(D), which inserts the entire document contacting the top-ranked sentence. For the instruction, we used the following template: "Context: <hint><context> Question: <question>".